OpenStudy (anonymous):

6x^4-13x^2+6=0 solve

OpenStudy (anonymous):

\[6(x ^{2})^{2}-(9+4)x ^{2}+6=0\] Can u proceed from here

OpenStudy (anonymous):

Put y = x^2. The equation becomes 6y^2 - 13y + 6 = 0 The discriminant is d = 169 - 144 > 0 so you have two real solutions for y y1 = (13 - 5)/12 = 2/3 and y2 = (13 + 5)/12 = 3/2 Since the original equation is of fourth degree, you need to have 4 solutions. So yo have to pose x1 = + sqrt(2/3) x2 = - sqrt(2/3) x3 = + sqrt(3/2) x4 = -sqrt(3/2) sqrt = square root